OpenStudy (anonymous):

What is the derivative of f(x) = 4^x?

OpenStudy (anonymous):

this is an exponential function

OpenStudy (anonymous):

I am not sure how you take the derivative of exponential function

OpenStudy (anonymous):

consider this: derivaitve (a^x)=a^x*ln(a)

OpenStudy (anonymous):

y= 4^x ln(y)= x ln 4 y'/y= ln 4 y' = ln 4 y y' = ln (4) 4^x

OpenStudy (anonymous):

This is it: f'(x)=4x^x-1